The nominal yield on 6-month T-bills is 4%, while default-free Japanese bonds that mature in 6 months have a nominal rate of 3%. In the spot exchange market, 1 yen equals $0.01. If interest rate parity holds, what is the 6-month forward exchange rate? Round the answer to five decimal places. Do not round intermediate calculations.
Answer #1

Interest Rate Parity Theory(IRPT) is given by

1+Rh / 1+Rf = F/S0

where

Rh - Risk-free interest rate in home country = 4/2 = 2%

Rf - Risk-free interest rate in foreign country = 3/2 = 1.5%

F - 6-month forward rate =?

S0 - Spot RAte =.01

Here exchange rate is given in the format of  dollar per yen. Hence, USA(first currency) is the home country and Japan(second currency) is the foreign country.

1.02/1.015 = F/.01

F/.01 = 1.00492610837

F = 1.00492610837*.01

= $0.0101 per yen

IRPT states that high interest rate in a country is offset by depreciation in the currency of that country. So here dollar should depreciate.

Here forward rate>spot rate. That is product in the exchange rate, yen is appreciating and dollar is depreciating, which is in line with IRPT.